c+45.6-38.25=112.76

Answer by usyim88hk(158)

This is a very simple equation
First, let us take a way the lousy numbers!!!
:
c+45.6-38.25 = 112.76
:
We can take away 38.25 by adding both sides by 38.25
:
c+45.6-38.25+38.25 = 112.76+38.25
c+45.6 = 151.01
:
Now take away 45.6 by subtracting 45.6 from both sides
:
c+45.6-45.6 = 151.01-45.6
c = 105.41
:
So c is equal to 105.41